# An interpretable family of projected normal distributions and a related copula model for Bayesian analysis of hypertoroidal data

### v1

@@ 5.1 MCMC steps for the TPN
5.2 General copula approach
Let $c$ be the density of the $d$-variate TPN with uniform margins ($TPN(\mathbf{0},\mathbf{0},\bm{\Sigma})$). If we want to use the proposed density as a copula model, with $j$-th marginal density equal to $f_{m,j}(\theta_{i,j})$ and CDF $F_{m,j}(\theta_{i,j})$, we have that the joint density of observation $i$ is equal to (20). Since $c$ has the functional form of a $TPN_{d}(\mathbf{0},\mathbf{0},\bm{\Sigma})$, it can be seen as the result of the integral
$c\left(\theta_{i,1}^{*},\dots,\theta_{i,d}^{*}\right)=\int_{(\mathbb{R}^{+})^{d}}f_{\bm{X}_{c,i}^{*}}(\bm{x}_{c,i}^{*})f_{\bm{X}_{s,i}^{*}}(\bm{x}_{s,i}^{*})\prod_{j=1}^{d}r_{i,j},$ (34)
where $f_{\bm{X}_{c,i}^{*}}(\bm{x}_{c,i}^{*})$ and $f_{\bm{X}_{s,i}^{*}}(\bm{x}_{s,i}^{*})$ in this case are the density of a $d$-variate normal with mean vector of zeros, covariance matrix $\bm{\Sigma}_{c}$ and $\bm{\Sigma}$ respectively, and
$\theta_{i,j}^{*}$ $=2\pi F_{m,j}(\theta_{i,j}-\mu_{j}),$ (35)
$\bm{x}_{c,i}^{*}$ $=\left(r_{i,1}\cos(\theta_{i,1}^{*}),\dots,r_{i,d}\cos(\theta_{i,d}^{*})\right),$ (36)
$\bm{x}_{s,i}^{*}$ $=\left(r_{i,1}\sin(\theta_{i,1}^{*}),\dots,r_{i,d}\sin(\theta_{i,d}^{*})\right).$ (37)
This means that if we introduce the latent variables $\mathbf{r}_{i}=(r_{i,1},\dots,r_{i,d})^{\top}$, we can work with the joint density
$f(\bm{\theta}_{i},\mathbf{r}_{i})=\left(f_{\bm{X}_{c,i}^{*}}(\bm{x}_{c,i}^{*})f_{\bm{X}_{s,i}^{*}}(\bm{x}_{s,i}^{*})\prod_{j=1}^{d}r_{i,j}\right)(2\pi)^{d}\prod_{j=1}^{d}f_{m,j}(\theta_{i,j}-\mu_{j}).$ (38)
Hence, the parameter $\bm{\Sigma}$ and the latent variables $\mathbf{r}_{i}$ can be sampled as in the previous section, whereas the variables $\bm{\theta}_{i}^{*}$, $\bm{x}_{c,i}^{*}$, and $\bm{x}_{s,i}^{*}$ must be used in place of $\bm{\theta}_{i}$, $\bm{x}_{c,i}$, and $\bm{x}_{s,i}$. The marginal density can depend on some parameters, that must be sampled using the Metropolis-Hastings algorithm, if closed-form full conditionals are not available.
@@ 6 Real Data Example
Figure 7: Posterior density of parameter $\lambda_{k,d}$ for the 6 circular variables. The y-axis is represented in logarithmic scale to better visualize the densitiess.
@@
To demonstrate the practical applicability of the proposed distribution, in this section, we apply it to a real dataset obtained from the National Data Buoy Center website (http://www.ndbc.noaa.gov/). We use data from four buoys located in the Gulf of Mexico, an ocean basin and a marginal sea of the Atlantic Ocean, off the Texas coast. Two buoys (Station PTAT2: 27.826° N, 97.051° W, and Station 42035: 29.235° N, 94.410° W) provide data only on wind direction, while the other two (Station 42019: 27.908° N, 95.343° W, and Station 42020: 26.970° N, 96.679° W) include both wind and wave directions: the spatial location of the buoys is shown in Figure 4. The data we use are daily measurements taken at 6 PM, from January 1st, 2020 to December 31st, 2020, yielding 366 recordings of a 6-dimensional vector of circular random variables. In the selected time window, station PTAT2 has 1 missing value. Station 42035 has 92 missing values, station 42019 has 157 missing values for wind direction and 57 for wave direction, while station 42020 has 46 missing values in wave direction and 51 in wind direction. The missing values are treated as additional random variables, and posterior values are simulated during the MCMC algorithm. A scatterplot for each pair of data is shown in Figure 5.
By using the proposed MCMC, we estimate the parameters of a 6 dimensional TPN as well as a copula model with wrapped Cauchy (WC) marginals. The WC density is formalized as
$f(\theta_{i,j}|\mu_{j}^{*},\lambda_{j})=\frac{1}{2\pi}\cdot\frac{1-\lambda_{j}^{2}}{1+\lambda_{j}^{2}-2\lambda_{j}\cos(\theta-\mu_{j}^{*})}$ (39)
and it depends on the mean direction $\mu_{j}^{*}$, which is also the mode of the distribution, and a concentration parameter $\lambda_{j}\in[0,1]$. Its CDF can also be written in closed form as
$F(\theta_{i,j}\leq a|\mu_{j}^{*},\lambda_{j})=\begin{cases}\frac{1}{2\pi}\arccos\left(\frac{(1+\lambda_{j}^{2})\cos(a-\mu_{j}^{*})-2\lambda_{j}}{1+\lambda_{j}^{2}-2\lambda_{j}\cos(a-\mu_{j}^{*})}\right),&\text{if }\sin(a-\mu_{j}^{*})\geq 0\\ 1-\frac{1}{2\pi}\arccos\left(\frac{(1+\lambda_{j}^{2})\cos(a-\mu_{j}^{*})-2\lambda_{j}}{1+\lambda_{j}^{2}-2\lambda_{j}\cos(a-\mu_{j})}\right),&\text{if }\sin(a-\mu_{j}^{*})<0\end{cases}$ (40)
where $a\in[0,2\pi)$. It should be notated that, by construction, the CDF at $\mu_{j}^{*}$ is equal to zero. Interestingly, it is possible to find in closed form also the quantile function of the WC which is
$F^{-1}(u|\mu_{j}^{*},\lambda_{j})=\left(\mu_{j}^{*}+\operatorname{sign}\left(0.5-u\right)\cdot\arccos\left(\frac{2\lambda_{j}+(1+\lambda_{j}^{2})\cos(2\pi u)}{1+\lambda_{j}^{2}+2\lambda_{j}\cos(2\pi u)}\right)\right).$ (41)
Equations (40) and (41), under the assumption $\mu_{j}^{*}=0$, can be used in (19). Since the WC is symmetrical, $\mu_{j}$ is the circular mean of the $j$-th marginal distribution.
We assume circular uniform prior distributions for the mean directions of both models, we assume $\kappa_{j}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,100000)I_{>0}$, $\lambda_{j}\sim U(0,1)$ and $\bm{\Sigma}\sim TIW(8,\mathbf{I}_{6})$. The MCMC is implemented with 30,000 iterations, with a burn-in of 10,000 and a thinning factor of 10. To select the best model between the two, we use the Continuous Ranked Probability Score (CRPS) (Gneiting and Raftery,, 2007). The CRPS is a metric that compares the true value of a random variable with its predictive distribution and is straightforward to compute in a Bayesian setting. In this real data example, it is computed using 10% of the data, which is set aside and not used in the model fitting. The CRPS for the TPN is 0.133 while for the WC is 0.117: since the WC is the one with lower CRPS index, results are only presented for this model.
In terms of both the mean direction $\mu_{k,d}$ (Figure 6) and the concentration parameter $\lambda_{k,d}$ (Figure 7), the distributions of the two wave variables appear quite similar, which is not the case for the wind directions. Figure 8 highlights a strong correlation between variables of the same type, while the correlation between wave and wind directions is weaker, though still around 0.5 for all combinations, even when recorded at the same buoy.
7 Conclusion
In this paper, we introduced two families of probability distributions designed for the Bayesian analysis of hypertoroidal data. The first family, the TPN distribution, is derived through the projection of multivariate normal distributions under specific parameter constraints. This construction ensures that the TPN distribution maintains key tractable properties, such as closure under marginalization and an intuitive parameter interpretation. The second family is the CTPN model, a copula-based extension of the TPN model, offering additional flexibility by allowing arbitrary univariate marginal distributions. Our proposal is the first copula model for vectors of any dimension of circular data using the projection approach.
The proposed distributions address some of the limitations of existing models, particularly by providing a clear interpretation of parameters governing the mean directions, concentration, and dependence structure. The use of the Markov Chain Monte Carlo (MCMC) framework for Bayesian estimation further enhances the applicability of these models, as demonstrated through simulations and real data applications.
Acknowledgments
@@
This work was partially supported by Mitsui Sumitomo Insurance Co., Ltd. (to Shogo Kato and Gianluca Mastrantonio) and JSPS KAKENHI Grant Number 25K15037 (to Shogo Kato).

### v2

@@ 5.1 MCMC steps for the TPN
5.2 The CTPN model
Let $c$ be the density of the $d$-variate TPN with uniform margins $\mathrm{TPN}_{d}(\bm{0},\bm{0},\bm{\Sigma})$. If we wish to use the proposed density as a copula model, with the $j$-th marginal density equal to $f_{j}(\theta_{i,j})$ and the corresponding marginal cumulative distribution function $F_{j}(\theta_{i,j})$, the joint density of observation $i$ is equal to (18). Since $c$ has the functional form of a $\mathrm{TPN}_{d}(\bm{0},\bm{0},\bm{\Sigma})$, it can be expressed as
$c\left(\theta_{i,1}^{*},\dots,\theta_{i,d}^{*}\right)=\int_{(\mathbb{R}^{+})^{d}}f_{\bm{X}_{c,i}^{*}}(\bm{x}_{c,i}^{*})f_{\bm{X}_{s,i}^{*}}(\bm{x}_{s,i}^{*})\left(\prod_{j=1}^{d}r_{i,j}\right)dr_{i,1}\cdots dr_{i,d},$ (30)
where $f_{\bm{X}_{c,i}^{*}}(\bm{x}_{c,i}^{*})$ and $f_{\bm{X}_{s,i}^{*}}(\bm{x}_{s,i}^{*})$ in this case denote the densities of $d$-variate normal with mean vector of zeros and covariance matrices $\bm{\Sigma}_{c}$ and $\bm{\Sigma}$, respectively, and
$\theta_{i,j}^{*}$ $=2\pi F_{j}(\theta_{i,j}-\mu_{j}),$ (31)
$\bm{x}_{c,i}^{*}$ $=\left(r_{i,1}\cos(\theta_{i,1}^{*}),\dots,r_{i,d}\cos(\theta_{i,d}^{*})\right),\quad\bm{x}_{s,i}^{*}=\left(r_{i,1}\sin(\theta_{i,1}^{*}),\dots,r_{i,d}\sin(\theta_{i,d}^{*})\right).$ (32)
This means that if we introduce the latent variables $\bm{r}_{i}=(r_{i,1},\dots,r_{i,d})^{\top}$, we can work with the joint density
$f(\bm{\theta}_{i},\bm{r}_{i})=\left(f_{\bm{X}_{c,i}^{*}}(\bm{x}_{c,i}^{*})f_{\bm{X}_{s,i}^{*}}(\bm{x}_{s,i}^{*})\prod_{j=1}^{d}r_{i,j}\right)(2\pi)^{d}\prod_{j=1}^{d}f_{j}(\theta_{i,j}-\mu_{j}).$ (33)
Hence, the parameter $\bm{\Sigma}$ and the latent variables $\bm{r}_{i}$ can be sampled as in the previous section, whereas the variables $\bm{\theta}_{i}^{*}$, $\bm{x}_{c,i}^{*}$, and $\bm{x}_{s,i}^{*}$ must be used in place of $\bm{\theta}_{i}$, $\bm{x}_{c,i}$, and $\bm{x}_{s,i}$. The marginal density can depend on some parameters, that must be sampled using the Metropolis-Hastings algorithm, if closed-form full conditionals are not available.
@@ 6 Real Data Example
To demonstrate the practical applicability of the proposed distribution, in this section, we apply it to a real dataset obtained from the National Data Buoy Center website (http://www.ndbc.noaa.gov/). We use data from four buoys located in the Gulf of Mexico, an ocean basin and a marginal sea of the Atlantic Ocean, off the Texas coast. Two buoys (Station PTAT2: 27.826° N, 97.051° W, and Station 42035: 29.235° N, 94.410° W) provide data only on wind direction, while the other two (Station 42019: 27.908° N, 95.343° W, and Station 42020: 26.970° N, 96.679° W) include both wind and wave directions: the spatial location of the buoys is shown in Fig. 4. The data we use are daily measurements taken at 6 PM, from January 1st, 2020 to December 31st, 2020, yielding 366 recordings of a 6-dimensional vector of circular random variables. In the selected time window, station PTAT2 has 1 missing value. Station 42035 has 92 missing values, station 42019 has 157 missing values for wind direction and 57 for wave direction, while station 42020 has 46 missing values in wave direction and 51 in wind direction. The missing values are treated as additional random variables, and posterior values are simulated during the MCMC algorithm. A scatterplot for each pair of data is shown in Fig. 5.
By using the proposed MCMC, we estimate the parameters of a 6 dimensional TPN as well as a CTPN with wrapped Cauchy marginals. The wrapped Cauchy density is defined as
$f_{j}(\theta_{i,j}|\mu_{j}^{*},\lambda_{j})=\frac{1}{2\pi}\cdot\frac{1-\lambda_{j}^{2}}{1+\lambda_{j}^{2}-2\lambda_{j}\cos(\theta_{i,j}-\mu_{j}^{*})}\quad(-\pi\leq\theta_{i,j}<\pi),$ (34)
where it depends on the location parameter $\mu_{j}^{*}$, which is also the mode of the distribution and its center of symmetry, and a concentration parameter $\lambda_{j}\in[0,1]$. Its cumulative distribution function can also be written in closed form as
$F_{j}(\theta_{i,j}\leq a|\mu_{j}^{*},\lambda_{j})=\begin{cases}\frac{1}{2\pi}\arccos\left\{\frac{(1+\lambda_{j}^{2})\cos(a-\mu_{j}^{*})-2\lambda_{j}}{1+\lambda_{j}^{2}-2\lambda_{j}\cos(a-\mu_{j}^{*})}\right\},&\text{if }\sin(a-\mu_{j}^{*})\geq 0\\ 1-\frac{1}{2\pi}\arccos\left\{\frac{(1+\lambda_{j}^{2})\cos(a-\mu_{j}^{*})-2\lambda_{j}}{1+\lambda_{j}^{2}-2\lambda_{j}\cos(a-\mu_{j})}\right\},&\text{if }\sin(a-\mu_{j}^{*})<0\end{cases}$ (35)
where $a\in[0,2\pi)$. By construction, the cumulative distribution function at $\mu_{j}^{*}$ is equal to zero. Interestingly, it is possible to find in closed form also the quantile function of the wrapped Cauchy which is
$F_{j}^{-1}(u|\mu_{j}^{*},\lambda_{j})=\mu_{j}^{*}+\operatorname{sign}\left(0.5-u\right)\cdot\arccos\left\{\frac{2\lambda_{j}+(1+\lambda_{j}^{2})\cos(2\pi u)}{1+\lambda_{j}^{2}+2\lambda_{j}\cos(2\pi u)}\right\}.$ (36)
Equations (35) and (36), under the assumption $\mu_{j}^{*}=0$, can be used in (17). Since the wrapped Cauchy is symmetrical, $\mu_{j}$ is the mean direction of the $j$-th marginal distribution.
We assume circular uniform prior distributions for the mean directions of both models, we assume $\kappa_{j}\stackrel{{\scriptstyle iid}}{{\sim}}{\cal N}_{1}(0,100000)I_{>0}$, $\lambda_{j}\sim U(0,1)$ and $\bm{\Sigma}\sim TIW(8,\bm{I}_{6})$. The MCMC is implemented with 30,000 iterations, with a burn-in of 10,000 and a thinning factor of 10. To select the best model between the two, we use the Continuous Ranked Probability Score (CRPS) (Gneiting2007). The CRPS is a metric that compares the true value of a random variable with its predictive distribution and is straightforward to compute in a Bayesian setting. In this real data example, it is computed using 10% of the data, which is set aside and not used in the model fitting. The CRPS for the TPN is 0.133 while for the wrapped Cauchy is 0.117: since the wrapped Cauchy is the one with lower CRPS index, results are only presented for this model.
@@
Figure 7: Posterior density of parameter $\lambda_{k,d}$ for the 6 circular variables.
@@
In terms of both the mean direction $\mu_{k,d}$ (Fig. 6) and the concentration parameter $\lambda_{k,d}$ (Fig. 7), the distributions of the two wave variables appear quite similar, which is not the case for the wind directions. Figure 8 highlights a strong correlation between variables of the same type, while the correlation between wave and wind directions is weaker, though still around 0.5 for all combinations, even when recorded at the same buoy. Overall, the three summaries in Figs. 68 convey a coherent picture. For wave directions, the posterior means $\mu_{k,d}$ and concentrations $\lambda_{k,d}$ are closely aligned across buoys, suggesting that the underlying generating mechanism is broadly shared across sites. By contrast, wind directions exhibit larger between-buoy variability in both location and concentration, which is consistent with local effects (e.g., orographic shielding, fetch) and differing measurement conditions. The dependence structure captured by $\bm{\Sigma}$ complements the marginal behavior: strong within-type correlations point to pronounced co-movement within the wave block and within the wind block, whereas cross-type correlations around 0.5 indicate a moderate but systematic coupling between waves and winds—even for co-located instruments. This pattern aligns with a physical interpretation in which wave fields integrate wind forcing over space and time, leading to smoother (and therefore more strongly correlated) dynamics for waves than for winds.
As a point of reference, more flexible hypertoroidal models can recover similar qualitative patterns but at the cost of reduced parameter interpretability. In our application, the proposed parametrization provides direct roles for $\mu_{k,d}$ (mean direction), $\lambda_{k,d}$ (concentration), and $\bm{\Sigma}$ (latent linear dependence), enabling clear, domain-relevant interpretations while capturing the main features of the data. A formal head-to-head comparison is beyond the scope of this section, but the qualitative agreement of locations, concentrations, and the sign/magnitude of cross-type dependence supports the practical interpretability of the proposed model.
Acknowledgement
@@
Shogo Kato and Gianluca Mastrantonio are grateful to Mitsui Sumitomo Insurance Co., Ltd. for financial support. The work of Kato is supported by JSPS KAKENHI Grant Number 25K15037.
Supplementary material
The Supplementary Material includes the proofs of the theoretical claims from the main article and simulation examples of the Bayesian estimation of the TPN and CTPN distributions.